Given the demand function P = 20 – 5Q, find the price elasticity of demand when price of the commodity is 5 Birr per unit. Mention if the demand is price elastic or inelastic at this point


1
Expert's answer
2021-08-17T16:58:46-0400

Solution:

Price elasticity of demand = "\frac{\triangle Q}{\triangle P} \times \frac{P}{Q}"

Demand function: P = 20 – 5Q

Derive the inverse demand function:

"Q = 4 - \frac{P}{5}"


"\frac{\triangle Q}{\triangle P} = -\frac{1}{5}"

P = 5

"Q = 4 - \frac{5}{5} = 4 - 1 = 3"

Price elasticity of demand = "- \frac{1}{5}\times \frac{5}{3}= -\frac{1}{3} = -0.33"

 

The demand is price inelastic since the value is less than one. This means that a change in price causes a smaller percentage change in demand or no effect at all.